# Discovering Large-Scale Structure at in the C3VO Survey

### v1

@@ 2.1.2 Spectroscopic Data
For our spectroscopic redshifts drawn from other surveys, we only adopted values that were denoted as secure by their reporting survey for our redshift range of interest (). Over half (50.5%) of these spectroscopic redshifts originated from VUDS (Le Fèvre et al., 2003), which primarily used photometric redshifts in the catalogs described above to select targets for spectroscopic follow-up with the Visible Multi-Object Spectrograph (VIMOS) on the Very Large Telescope (VLT). Details regarding the survey design, observations, data reduction, and redshift determination can be found in Le Fèvre et al. (2015), and a full description of the first VUDS data release can be found in Tasca et al. (2017). We consider galaxies with flags X2, X3, X4, and X9 as secure, where X may equal 0, 1, 2, or 3, according to the flagging code employed in Le Fèvre et al. (2015). Such flags indicate a redshift that is reliable at the 75 to 99.3% level.
The next largest source (24.5%) of our spectroscopic redshift sample came from the Bright and Deep phases of the zCOSMOS survey (Lilly et al., 2007, 2009; Diener et al., 2013, 2015). Other spectroscopic redshifts in the COSMOS field were taken from Casey et al. (2015); Chiang et al. (2015); Diener et al. (2015) at . The 3D-HST, VANDELS, VVDS, and ACES surveys (including Le Fèvre et al. 2004, 2013; Vanzella et al. 2009; Hathi et al. 2009; Straughn et al. 2009; Balestra et al. 2010; Cooper et al. 2012; Kurk et al. 2013; Trump et al. 2013; Morris et al. 2015; McLure et al. 2018; Pentericci et al. 2018; Talia et al. 2023 among others) that cover the ECDFS field make up 17.0% of the spectroscopic redshift sample (N. Hathi, priv. comm.). The field galaxies in the CFHTLS-D1 field are sourced from the Deep and UltraDeep phases of the VIMOS VLT Deep Survey (VVDS; Le Fèvre et al., 2005, 2013) which make up 5.5% of the spectroscopic redshift sample. Although each spectroscopic survey has its own unique flagging system, we chose galaxies with flags as close as possible to what would be considered secure in VUDS. If a galaxy appeared in multiple surveys, we adopted the redshift with the highest confidence flag, with the VUDS redshift preferred in the case of a tie.
@@ 2.2 C3VO Observations
The remainder of the redshifts were drawn from dedicated C3VO spectroscopy, which was obtained with the Multi-Object Spectrometer For Infra-Red Exploration (MOSFIRE; McLean et al., 2012) and the DEep Imaging Multi-Object Spectrograph (DEIMOS; Faber et al., 2003) on the Keck I/II telescopes. The C3VO survey’s Keck observations were designed with the goal of a near complete magnitude-limited survey to of the six most significant overdensities found in VUDS, including those reported in Lemaux et al. (2014b, 2018); Cucciati et al. (2014, 2018); Shen et al. (2021); Forrest et al. (2023); Shah et al. (2024a); Staab et al. (2024). The observations targeted star-forming galaxies of all types to a magnitude of , magnitudes at which we could hope to recover continuum redshifts, and fainter sources () for which we would only be able to spectroscopically confirm Lyman- emitting galaxies. Details of the C3VO observations of all structures can be found in the references given above.
@@
For this work, we do not incorporate all C3VO observations, but rather only those from select overdensities as our goal is primarily to assess our ability to recover structure from broad-breadth spectroscopic surveys rather than those aimed at dedicated followup. For consistency with some of the previous C3VO works, we include the initial DEIMOS followup observations of the Taralay protostructure at in the COSMOS field as described in Lemaux et al. (2022) and the MOSFIRE observations of PClJ0227-0421 at in the CFHTLS-D1 field used in Shen et al. (2021). Both of these observations add considerably to the spectroscopic sampling at the high-redshift end of their respective fields. Additionally, we incorporate spectroscopic observations of the ECDFS field described in Shah et al. (2024a), which includes both C3VO observations and a newly compiled catalog of public redshifts made by one of the co-authors (NPH). As can be seen in Figure 1 of Shah et al. (2024a), the former comprise only a small fraction of the redshifts in the field, and are primarily targeted at the Smruti protostructure at . The details of these observations and the reduction of these data can be found in the corresponding reference listed above as well as Forrest et al. (2023, 2024).
@@
Spectroscopic redshifts from DEIMOS data were determined with the publicly available Deep Evolutionary Extragalactic Probe 2 (DEEP2; Davis et al., 2003; Newman et al., 2013) redshift measurement program zspec (Newman et al., 2013). Since this program was intended primarily to evaluate the spectral redshift of galaxies at , empirical high-redshift galaxy templates from VUDS and VVDS were added to the base template set in the program, as were high-resolution empirical Lyman- templates from Lemaux et al. (2009). All observed objects targeted or otherwise were visually inspected and assigned a spectroscopic redshift and a redshift quality code . Under the DEEP2 convention, secure stellar ( = -1) and extragalactic ( = 3, 4) redshifts are scientifically usable at the 95% confidence level (Newman et al., 2013).
@@
MOSFIRE redshifts were initially determined by applying an identical methodology to the one described above for the DEIMOS spectra following a visual inspection for intermediate to high signal-to-noise (S/N) emission features. A second pass redshift estimation was performed after a more comprehensive visual inspection of all two-dimensional MOSFIRE spectra, both to identify lower S/N features and detections of non-targets that serendipitously fell in the MOSFIRE slits. A one-dimensional extraction was performed at the spatial location of emission lines identified during the visual inspection process and redshifts were determined via model fitting (see Forrest et al. 2023 for more details). Redshift flags for MOSFIRE spectra followed the convention of those employed for the DEIMOS redshift determination.
@@
In addition to the high- C3VO redshifts, we also incorporated Keck/DEIMOS observations in the CFHTLS-D1 field of the cluster XLSS005 from the ORELSE survey (see Lemaux et al. 2019 and references therein for details). This campaign yielded 750 secure spectroscopic redshifts. While these redshifts are exclusively at , they are useful as a means of eliminating lower-redshift interlopers in our galaxy density mapping (see §3.2).
@@
Combining all data sets and imposing the NIR apparent magnitude cuts used to define both our photometric and spectroscopic sample, as discussed in the next section, resulted in a total of 7451 unique galaxies with a high-quality in the redshift range 333For C3VO observations “high-quality” spectral redshifts refers only to flags -1, 3, & 4. The VVDS, zCOSMOS, and our public compilation of redshifts are placed on the VUDS flagging system, where the term high-quality spectral redshifts refers to all subclasses of flags 2, 3, 4, & 9 (see Lemaux et al. 2022 for more details on the reliability of these flags)..
@@ 3.1 Photometric and Spectroscopic Redshifts
Applying the photometric cuts detailed above to both the spectroscopic and photometric sample, the spectroscopic redshift fraction444This quantity, which is abbreviated later in the paper as SF, refers to the fraction of objects brighter than the photometric limit imposed in a given field in the range with a secure spectral redshift. ranges from 0.06 to 0.2 across the three fields, with CFHTLS-D1 having the sparsest coverage and ECDFS having the densest. For the COSMOS, ECDFS, and CFHTLS-D1 fields, the scatter is 0.013, 0.017, and 0.043, with catastrophic outlier rates of 8.8%, 12.0%, and 14%, and biases (i.e., (-)/(1+)) of -0.0019, 0.0009, and -0.0149, respectively, as measured for galaxies in the redshift range subject to the same photometric cuts as above.
3.2 Voronoi Tessellation Monte-Carlo Mapping
@@
Hung et al. (2020) successfully used a powerful technique known as Voronoi tessellation Monte-Carlo (VMC) mapping to find hundreds of candidate galaxy clusters and groups over a combined spectroscopic footprint of 1.4 deg2 in the ORELSE survey at . Despite the increase in data quality from the ORELSE survey in terms of overall number of spectroscopic redshifts, as well as the depth and breadth of the imaging data, we expect that the same structure search algorithm will have a more challenging time producing results that are as fruitful at higher redshift. This expectation is due partially to the lower fraction of spectroscopy available, as the number counts of galaxies increase precipitously with deeper imaging, as well as the larger photometric redshift errors that accompany higher redshift galaxies. Additionally, the density contrast of forming structures at high redshift is less than that of massive clusters, and more resemble the density contrast of groups (see, e.g., Forrest et al. 2024). Details on the construction of the VMC maps for both and at can be found in Tomczak et al. (2017); Lemaux et al. (2017, 2018); Hung et al. (2020); Lemaux et al. (2022), but we will summarize the process here. For this work, we adopt the particular approach of Lemaux et al. (2022).
@@
Voronoi tessellation has broad applications that also include being capable of acting as a type of density field estimator. It defines a polygonal cell for every object in a 2D plane where the area inside the cell is closer to the host object than any other object. The cell size is thus indicative of the density at a given location; the smaller the cell size, the higher the density. The survey data cover a broad range in redshift, the data are cut into redshift slices, each 7.5 Mpc in width, or 0.015–0.08 for 2–5. We choose this approach rather than a three-dimensional tesselation primarily because induced motion complicates the interpretation of line of sight positions. By breaking the data up into two-dimensional slices of this size we are able to marginalize over velocity substructure in forming structures. A tessellation is performed on each slice and the resultant density field is sampled onto a grid of 75 square kpc pixels. In order to minimize the chances of splitting individual structures across slices, we define the slices such that two sequential slices will have 90% overlap in redshift.
@@
In every redshift slice, there are galaxies with spectroscopic and photometric redshifts. Photometric redshifts have much larger uncertainties than spectroscopic redshifts, and we account for this through a Monte-Carlo method. For each Monte-Carlo realization, we sample the reconstructed PDF of every using an asymmetric Gaussian, which will shift some photometric redshifts in or out of the redshift boundaries of a particular slice and accordingly change the measured density values. Spectroscopic redshifts are also treated statistically according to their confidence flags, with reliability thresholds adopted from those reported in Lemaux et al. (2022). We then perform the Voronoi tessellation on all of the and objects that fall in given slice for a given iteration. This process is run a total of 100 times for each slice, and the final VMC density map for that slice is obtained by median combining the densities from all realizations. We can then obtain the local overdensity in a pixel () with
@@ 3.3 Detecting Coherent Structure
With the VMC maps in hand in each field over the redshift range , we apply our structure search algorithm. Significant detections in each slice of the VMC maps, which represent potential signal from a structure, are found using the standard photometry software package Source Extractor (SExtractor; Bertin & Arnouts, 1996). All slices of the VMC maps in each field are fed into SExtractor and the outer 20-30% of the area of the maps in each redshift slice is masked to excise areas of the map where no data exist and to mitigate edge effects. The median density of each slice as measured in the inner 70-80% of the maps is used as a background level. Unlike previous works (e.g., Hung et al., 2020), we do not employ a fit of the median density as a function of redshift to determine the background. This choice was made to mitigate issues with artificial clustering in redshift space of that spuriously cause wide peaks or troughs in the density distribution (Fig. 1), which lead to the suppression of of real structure and enhancement of false positives if a fit is instead employed.
@@
The root mean squared fluctuations in the density field (RMS), which is also measured in the inner 70-80% of the maps, varies much more rapidly and is much more sensitive to the presence of real structure (Fig. 2). As a result, we employ an fifth-order polynomial fit, iteratively clipped at , of the RMS as a function of redshift for each field and adopt the fitted values per slice when running SExtractor on the VMC maps. In most cases, the structures only occupy a small fraction of the VMC map’s area. The iterative clipping is most beneficial for the ECDFS field, where the structures can be 30 to 40% of the area of the map, leading to increases in the average local density and the RMS driven by astrophysical reasons. Note that the RMS measured in SExtractor is different than a traditional measure of the RMS of all of the pixels in a given slice. Rather, the SExtractor RMS is measured on the iteratively 3-clipped distribution of the average density values in regions whose sizes are dictated by the BACK_SIZE input parameter (64x64 pixels in our case). Once determined for each slice, the median density and RMS are then used by SExtractor, along with several other parameters, to detect signal from potential structure in each density slice.
@@
Four SExtractor parameters in particular are especially pertinent for this work. The DETECT_THRESH and DETECT_MINAREA parameters define the base thresholds for what constitutes a detection in SExtractor. DETECT_THRESH sets how much higher the density floor must be relative to the RMS noise in the background. DETECT_MINAREA sets the minimum number of contiguous pixels. The detection parameters are best chosen to maximize completeness while also minimizing spurious detections due to noise. Tests with the ORELSE data have suggested that DETECT_THRESH values of around 4 is a reasonable choice to balance maximizing the purity and completeness of detections (Hung et al., 2020). While our overall spectroscopic catalog is larger with the C3VO data, the fraction of spectroscopic redshifts is generally lower, primarily due to the extreme faintness of the galaxy population that is adopted for our density mapping (IRAC1 24.8). In addition, the density contrast of forming clusters (i.e., protoclusters) is considerably less than the cores of established clusters by an order of magnitude or more even in the core regions (see, e.g., Forrest et al. 2024). As such, in this work we adopt a less restrictive threshold of 3. Additionally, we adopt a DETECT_MINAREA value of 100 pixels. Such an area is equivalent to 0.5625 Mpc2 (corresponding to, e.g., 9 comoving Mpc2 at ). This is roughly two times smaller than the sizes of lower mass protoclusters (i.e., M⊙) at these redshifts (e.g., Chiang et al., 2013; Lim et al., 2024), which allows us to probe down to the protogroup scale.
@@
The DEBLEND_NTHRESH and DEBLEND_MINCONT parameters define the deblending parameters which can split single contiguous detections into multiple pieces. DEBLEND_NTHRESH sets the number of deblending sub-thresholds which are exponentially spaced between the detection floor and the peak of the detection. DEBLEND_MINCONT sets the minimum contrast which is the fraction of the overdensity in a substructure relative to the overdensity in the entire structure needed to be considered a separate detection. Deblending is desirable if there are blended systems in a field that may consist of two or more sub-structures, but setting the deblending too high can split single structures. As we are more interested in detecting structure than in resolving the individual sub-components of structures, we adopt a less aggressive deblending scheme which uses a DEBLEND_NTHRESH value of 12 and DEBLEND_MINCONT value of 0.3. Such a choice will only break up the most conspicuous of multicomponent systems.
@@
Once the individual detections in each redshift slice are identified with SExtractor, we link the detections across redshift slices. SExtractor reports the density-weighted barycenter (hereafter simply barycenter) in right ascension and declination (hereafter RA and Dec, respectively) for every detection. We can check if a pair of detections across two sequential slices are part of the same chain by calculating the transverse distance between their barycenters. If they are within a linking radius, which we set to 2 Mpc after testing values between 1.5 and 3 Mpc in 0.5 Mpc increments, we combine the two detections into one linked chain and obtain a new barycenter from the positions of the individual detections weighted by their isophotal flux as defined by SExtractor. It is important to note that because we applied SExtractor to density maps, the flux values are not traditional photometric values, but rather represent the level of galaxy overdensity. The linking continues until no more detections are found that can be added to the chain.
@@
We find all possible linked chains starting from the lowest redshift slice. Consequently, we will have many linked chains that are either wholly subsets of larger linked chains that were started at lower redshifts or differ by a few detections in the middle of the chain. In order to avoid potentially double counting the same protostructure candidate, we first remove any linked chain that is a complete subset of another chain. We obtain the protostructure candidates by fitting a Gaussian to the isophotal fluxes of the linked detections as a function of redshift.
@@
In some cases, the linked chain extends over very wide redshift ranges () and contain multiple peaks. These long linked chains contain about 10% of all peaks we identify. We identify peaks in a linked chain by first sorting all isophotal flux values in the chain, taking note of the redshift of the highest flux point. We then move on to the next highest flux point and examine its redshift. If it is within of the first flux point, we associate the two points together into the same peak. This threshold was chosen as it is the expected line-of-sight extent of a protocluster plus induced motion, i.e., 10 to 30 comoving Mpc (Chiang et al., 2013; Lim et al., 2024). Otherwise, the point is categorized as a second peak. We continue this from high to low flux values, checking whether each point is within the minimum and maximum redshift range of each identified peak and associating the points accordingly, until we reach 20% of the original highest flux point of the linked chain. We set this threshold in order to reduce complications with noise around the flux floor of the linked chain.

### v2

@@ 2.1.2 Spectroscopic Data
For our spectroscopic redshifts drawn from other surveys, we only adopted values that were denoted as secure by their reporting survey for our redshift range of interest (). Over half (50.5%) of these spectroscopic redshifts originated from VUDS (Le Fèvre et al., 2003), which primarily used photometric redshifts in the catalogs described above to select targets for spectroscopic follow-up with the VIMOS on the Very Large Telescope (VLT). Details regarding the survey design, observations, data reduction, and redshift determination can be found in Le Fèvre et al. (2015), and a full description of the first VUDS data release can be found in Tasca et al. (2017). We consider galaxies with flags X2, X3, X4, and X9 as secure, where X may equal 0, 1, 2, or 3, according to the flagging code employed in Le Fèvre et al. (2015). Such flags indicate a redshift that is reliable at the 75–99.3% level.
The next largest source (24.5%) of our spectroscopic redshift sample came from the Bright and Deep phases of the zCOSMOS survey (Lilly et al., 2007, 2009; Diener et al., 2013, 2015). Other spectroscopic redshifts in the COSMOS field were taken from Casey et al. (2015), Chiang et al. (2015), and Diener et al. (2015) at . The 3D-HST, VANDELS, VVDS, and ACES surveys (including Le Fèvre et al. 2004, 2013; Vanzella et al. 2009; Hathi et al. 2009; Straughn et al. 2009; Balestra et al. 2010; Cooper et al. 2012; Kurk et al. 2013; Trump et al. 2013; Morris et al. 2015; McLure et al. 2018; Pentericci et al. 2018; Talia et al. 2023 among others) that cover the ECDFS field make up 17.0% of the spectroscopic redshift sample (N. Hathi 2025, private communication). The field galaxies in the CFHTLS-D1 field are sourced from the Deep and UltraDeep phases of the VIMOS VLT Deep Survey (VVDS; Le Fèvre et al., 2005, 2013), which make up 5.5% of the spectroscopic redshift sample. Although each spectroscopic survey has its own unique flagging system, we chose galaxies with flags as close as possible to what would be considered secure in VUDS. If a galaxy appeared in multiple surveys, we adopted the redshift with the highest confidence flag, with the VUDS redshift preferred in the case of a tie.
@@ 2.2 C3VO Observations
The remainder of the redshifts were drawn from dedicated C3VO spectroscopy, which was obtained with the Multi-Object Spectrometer For Infra-Red Exploration (MOSFIRE; McLean et al., 2012) and the DEep Imaging Multi-Object Spectrograph (DEIMOS; Faber et al., 2003) on the Keck I/II telescopes. The C3VO survey’s Keck observations were designed with the goal of a near complete magnitude-limited survey to of the six most significant overdensities found in VUDS, including those reported in Lemaux et al. (2014b, 2018), Cucciati et al. (2014, 2018), Shen et al. (2021), Forrest et al. (2023), Shah et al. (2024a), and Staab et al. (2024). The observations targeted star-forming galaxies of all types to a magnitude of , magnitudes at which we could hope to recover continuum redshifts, and fainter sources () for which we would only be able to spectroscopically confirm Ly emitting galaxies. Details of the C3VO observations of all structures can be found in the references given above.
@@
For this work, we do not incorporate all C3VO observations, but rather only those from select overdensities, as our goal is primarily to assess our ability to recover structure from broad-breadth spectroscopic surveys rather than those aimed at dedicated follow-up. For consistency with some of the previous C3VO works, we include the initial DEIMOS follow-up observations of the Taralay protostructure at in the COSMOS field as described in Lemaux et al. (2022) and the MOSFIRE observations of PClJ0227-0421 at in the CFHTLS-D1 field used in Shen et al. (2021). Both of these observations add considerably to the spectroscopic sampling at the high-redshift end of their respective fields. Additionally, we incorporate spectroscopic observations of the ECDFS field described in Shah et al. (2024a), which includes both C3VO observations and a newly compiled catalog of public redshifts made by one of the coauthors (N.P.H.). As can be seen in Figure 1 of Shah et al. (2024a), the former comprise only a small fraction of the redshifts in the field, and are primarily targeted at the Smruti protostructure at . The details of these observations and the reduction of these data can be found in the corresponding reference listed above as well as Forrest et al. (2023, 2024).
@@
Spectroscopic redshifts from DEIMOS data were determined with the publicly available Deep Evolutionary Extragalactic Probe 2 (DEEP2; Davis et al., 2003; Newman et al., 2013) redshift measurement program zspec (Newman et al., 2013). Since this program was intended primarily to evaluate the spectral redshift of galaxies at , empirical high-redshift galaxy templates from VUDS and VVDS were added to the base template set in the program, as were high-resolution empirical Ly templates from Lemaux et al. (2009). All observed objects targeted or otherwise were visually inspected and assigned a spectroscopic redshift and a redshift quality code . Under the DEEP2 convention, secure stellar () and extragalactic (, 4) redshifts are scientifically usable at the 95% confidence level (Newman et al., 2013).
@@
MOSFIRE redshifts were initially determined by applying an identical methodology to the one described above for the DEIMOS spectra following a visual inspection for intermediate to high signal-to-noise ratio (S/N) emission features. A second pass redshift estimation was performed after a more comprehensive visual inspection of all two-dimensional MOSFIRE spectra, both to identify lower S/N features and detections of non-targets that serendipitously fell in the MOSFIRE slits. A one-dimensional extraction was performed at the spatial location of emission lines identified during the visual inspection process, and redshifts were determined via model fitting (see Forrest et al. 2023 for more details). Redshift flags for MOSFIRE spectra followed the convention of those employed for the DEIMOS redshift determination.
@@
In addition to the high- C3VO redshifts, we also incorporated Keck/DEIMOS observations in the CFHTLS-D1 field of the cluster XLSS005 from the ORELSE survey (see Lemaux et al. 2019 and references therein for details). This campaign yielded 750 secure spectroscopic redshifts. While these redshifts are exclusively at , they are useful as a means of eliminating lower-redshift interlopers in our galaxy density mapping (see Section 3.2).
@@
Combining all data sets and imposing the NIR apparent magnitude cuts used to define both our photometric and spectroscopic sample, as discussed in Section 3, resulted in a total of 7451 unique galaxies with a high-quality in the redshift range 333For C3VO observations “high-quality” spectral redshifts refers only to flags -1, 3, & 4. The VVDS, zCOSMOS, and our public compilation of redshifts are placed on the VUDS flagging system, where the term high-quality spectral redshifts refers to all subclasses of flags 2, 3, 4, & 9 (see Lemaux et al. 2022 for more details on the reliability of these flags)..
@@ 3.1 Photometric and Spectroscopic Redshifts
Applying the photometric cuts detailed above to both the spectroscopic and photometric sample, the global spectroscopic redshift fraction444The quantity here refers to the fraction of objects brighter than the photometric limit imposed in a given field in the range with a secure spectral redshift. In contrast, the “spectroscopic redshift fraction,” which is introduced later in the paper and is abbreviated as “SF,” refers to a multiplicative factor of the fraction of galaxies with a secure spectroscopic redshifts in different photo- and magnitude bins in the COSMOS field, i.e., a scaling up or down of our spectroscopic selection function in the COSMOS field (see Section 4.2.) ranges from 0.06–0.2 across the three fields, with CFHTLS-D1 having the sparsest coverage and ECDFS having the densest. For the COSMOS, ECDFS, and CFHTLS-D1 fields, the scatter is 0.013, 0.017, and 0.043, with catastrophic outlier rates of 8.8%, 12.0%, and 14%, and biases (i.e., (-)/(1+)) of -0.0019, 0.0009, and -0.0149, respectively, as measured for galaxies in the redshift range subject to the same photometric cuts as above.
3.2 Voronoi Tessellation Monte Carlo Mapping
@@
Hung et al. (2020) successfully used a powerful technique known as Voronoi tessellation Monte Carlo (VMC) mapping to find hundreds of candidate galaxy clusters and groups over a combined spectroscopic footprint of 1.4 deg2 in the ORELSE survey at . Despite the increase in data quality from the ORELSE survey in terms of overall number of spectroscopic redshifts, as well as the depth and breadth of the imaging data, we expect that the same structure search algorithm will have a more challenging time producing results that are as fruitful at higher redshift. This expectation is due partially to the lower fraction of spectroscopy available, as the number counts of galaxies increase precipitously with deeper imaging, as well as the larger photometric redshift errors that accompany higher redshift galaxies. Additionally, the density contrast of forming structures at high redshift is less than that of massive clusters, and more resembles the density contrast of groups (see, e.g., Forrest et al. 2024). Details on the construction of the VMC maps for both and at can be found in Tomczak et al. (2017); Lemaux et al. (2017, 2018); Hung et al. (2020); Lemaux et al. (2022), but we will summarize the process here. For this work, we adopt the particular approach of Lemaux et al. (2022).
@@
Voronoi tessellation has broad applications that also include being capable of acting as a type of density field estimator. It defines a polygonal cell for every object in a two-dimensional plane where the area inside the cell is closer to the host object than any other object. The cell size is thus indicative of the density at a given location; the smaller the cell size, the higher the density. The survey data cover a broad range in redshift, and the data are cut into redshift slices, each 7.5 Mpc in width, or 0.015–0.08 for 2–5. We choose this approach rather than a three-dimensional tessellation primarily because peculiar velocities complicate the interpretation of line-of-sight positions. By breaking the data up into two-dimensional slices of this size, we are able to marginalize over velocity substructure in forming structures. A tessellation is performed on each slice and the resultant density field is sampled onto a grid of 75 square kpc pixels. In order to minimize the chances of splitting individual structures across slices, we define the slices such that two sequential slices will have 90% overlap in redshift.
@@
In every redshift slice, there are galaxies with spectroscopic and photometric redshifts. Photometric redshifts have much larger uncertainties than spectroscopic redshifts, and we account for this through a Monte Carlo method. For each Monte Carlo realization, we sample the reconstructed probability density function (PDF) of every using an asymmetric Gaussian, which will shift some photometric redshifts in or out of the redshift boundaries of a particular slice and accordingly change the measured density values. Spectroscopic redshifts are also treated statistically according to their confidence flags, with reliability thresholds adopted from those reported in Lemaux et al. (2022). We then perform the Voronoi tessellation on all of the and objects that fall in a given slice for a given realization. The Gaussian sampling on all photometric redshifts and subsequent Voronoi tessellation is run a total of 100 times for each slice, and the final VMC density map for that slice is obtained by median combining the densities from all realizations. We can then obtain the local overdensity in a pixel () with
@@ 3.3 Detecting Coherent Structure
With the VMC maps in hand, in each field over the redshift range , we apply our structure search algorithm. Significant detections in each slice of the VMC maps, which represent potential signal from a structure, are found using the standard photometry software package Source Extractor (SExtractor; Bertin & Arnouts, 1996). All slices of the VMC maps in each field are fed into an initial run of SExtractor and the outer 20–30% of the area of the maps in each redshift slice is masked to excise areas of the map where no data exist and to mitigate edge effects. In this initial run, median density of each slice as measured in the inner 70–80% of the maps and is used as a background level. This run is also used to set the rms fluctuations in the density field (see below), which, along with the median density, is used to set the detection threshold for structure in the maps. Unlike previous works (e.g., Hung et al., 2020), we do not employ a fit of the median density as a function of redshift to determine the background. This choice was made to mitigate issues with artificial clustering in redshift space of that spuriously cause wide peaks or troughs in the density distribution (Fig. 1), which lead to the suppression of of real structure and enhancement of false positives if a fit is instead employed.
@@
The rms of the density field, which is also measured in the inner 70-80% of the maps, varies much more rapidly and is much more sensitive to the presence of real structure (Fig. 2). As a result, we employ a fifth-order polynomial fit, iteratively clipped at , of the rms as a function of redshift for each field and adopt the fitted values per slice when running SExtractor on the VMC maps. In most cases, the structures only occupy a small fraction of the VMC map’s area. The iterative clipping is most beneficial for the ECDFS field, where the structures can be 30–40% of the area of the map, leading to increases in the average local density and the rms driven by astrophysical reasons. Note that the rms measured in SExtractor is different than a traditional measure of the rms of all of the pixels in a given slice. Rather, the SExtractor rms is measured on the iteratively 3-clipped distribution of the average density values in regions whose sizes are dictated by the BACK_SIZE input parameter (6464 pixels in our case). Once determined for each slice, the median density and rms are then used by SExtractor, along with several other parameters, to detect the signal from potential structure in each density slice. Note that in this final run of SExtractor, we no longer apply masking, but rather identify those candidates in our final protostructure candidate catalog whose overdensity-weighted barycenters lie within 2 Mpc of the masked region in the maps, as their authenticity is likely more suspect and the strength of the detection is possibly affected by their proximity to the edges of the maps.
@@
Four SExtractor parameters in particular are especially pertinent for this work. The DETECT_THRESH and DETECT_MINAREA parameters define the base thresholds for what constitutes a detection in SExtractor. DETECT_THRESH sets how much higher the density floor must be relative to the rms noise in the background. DETECT_MINAREA sets the minimum number of contiguous pixels. The detection parameters are best chosen to maximize completeness while also minimizing spurious detections due to noise. Tests with the ORELSE data have suggested that DETECT_THRESH values of around 4 is a reasonable choice to balance maximizing the purity and completeness of detections (Hung et al., 2020). While our overall spectroscopic catalog is larger with the C3VO data, the global fraction of spectroscopic redshifts is generally lower, primarily due to the extreme faintness of the galaxy population that is adopted for our density mapping (IRAC1 24.8). In addition, the density contrast of forming clusters (i.e., protoclusters) is considerably less than the cores of established clusters by an order of magnitude or more even in the core regions (see, e.g., Forrest et al. 2024). As such, in this work we adopt a less restrictive threshold of 3. Additionally, we adopt a DETECT_MINAREA value of 100 pixels. Such an area is equivalent to 0.5625 Mpc2 (corresponding to, e.g., 9 comoving Mpc2 at ). This is roughly two times smaller than the sizes of lower-mass protoclusters (i.e., M⊙) at these redshifts (e.g., Chiang et al., 2013; Lim et al., 2024), which allows us to probe down to the protogroup scale.
@@
The DEBLEND_NTHRESH and DEBLEND_MINCONT parameters define the deblending parameters that can split single contiguous detections into multiple pieces. DEBLEND_NTHRESH sets the number of deblending subthresholds that are exponentially spaced between the detection floor and the peak of the detection. DEBLEND_MINCONT sets the minimum contrast, which is the fraction of the overdensity in a substructure relative to the overdensity in the entire structure needed to be considered a separate detection. Deblending is desirable if there are blended systems in a field that may consist of two or more substructures, but setting the deblending too high can split single structures. As we are more interested in detecting structure than in resolving the individual subcomponents of structures, we adopt a less aggressive deblending scheme that uses a DEBLEND_NTHRESH value of 12 and DEBLEND_MINCONT value of 0.3. Such a choice will only break up the most conspicuous of multicomponent systems, though in practice we had no cases of deblending with this choice of parameters.
If the deblending were more aggressive, we could more effectively separate the subcomponents of more complex structures, such as Hyperion, Elentári, Smruti, Ruchi, and PClJ0227–0421 to take a few examples from the literature. However, by effectively turning off the deblending, we can ensure greater consistency of detections between redshift slices rather than risk having subcomponents deblended in some slices but not in neighboring slices. Through this, we maximize our chances of detecting the overall structure for a given system, but as we will necessarily blur out subcomponents of complex systems, we expect a poorer performance in our detection algorithm’s ability to detect all subcomponents of complex structures known from the literature (see Section 5.2).
@@
Once the individual detections in each redshift slice are identified with SExtractor, we link the detections across redshift slices. SExtractor reports the density-weighted barycenter (hereafter simply barycenter) in R.A. and decl. for every detection. We can check if a pair of detections across two sequential slices are part of the same chain by calculating the transverse distance between their barycenters. If they are within a linking radius, which we set to 2 Mpc (chosen to be slightly longer than the optimal length found through tests with the ORELSE data in Hung et al. 2020 to account for the greater sizes of protoclusters relative to clusters), we combine the two detections into one linked chain and obtain a new barycenter from the positions of the individual detections weighted by their isophotal flux as defined by SExtractor. It is important to note that because we applied SExtractor to density maps, the flux values are not traditional photometric values, but rather represent the level of galaxy overdensity. The linking continues until no more detections are found that can be added to the chain.
@@
We find all possible linked chains starting from the lowest-redshift slice. Consequently, we will have many linked chains that are either wholly subsets of larger linked chains that were started at lower redshifts or differ by a few detections in the middle of the chain. In order to avoid potentially double-counting the same protostructure candidate, we first remove any linked chain that is a complete subset of another chain. We obtain the protostructure candidates by fitting a Gaussian to the isophotal fluxes of the linked detections as a function of redshift.
@@
In some cases, the linked chain extends over very wide redshift ranges () and contain multiple peaks. These long linked chains contain about 10% of all peaks we identify. We identify peaks in a linked chain by first sorting all isophotal flux values in the chain, taking note of the redshift of the highest flux point. We then move on to the next highest flux point and examine its redshift. If it is within of the first flux point, we associate the two points together into the same peak. This threshold was chosen as it is the expected line-of-sight extent of a protocluster plus induced motion, i.e., 10–30 comoving Mpc (Chiang et al., 2013; Lim et al., 2024). Otherwise, the point is categorized as a second peak. We continue this from high to low flux values, checking whether each point is within the minimum and maximum redshift range of each identified peak and associating the points accordingly, until we reach 20% of the original highest flux point of the linked chain. We set this threshold in order to reduce complications with noise around the flux floor of the linked chain.